OpenStudy (anonymous):

sinØ= cos(90-Ø) i said its false but im nt sure

OpenStudy (anonymous):

That is true. :)

OpenStudy (anonymous):

its not false?

OpenStudy (primeralph):

it is not false.

OpenStudy (anonymous):

thank youuu :)

OpenStudy (primeralph):



OpenStudy (primeralph):

nothing for me?

OpenStudy (anonymous):

i just have one question the 0 i think its called thetha i dnt get the concept of that :/

OpenStudy (primeralph):

it is called theta

OpenStudy (anonymous):

@primeralph lol what do u want

OpenStudy (primeralph):

@nisa never mind.

OpenStudy (anonymous):

lol i wish they d let us give more medals

OpenStudy (ajprincess):

\(\cos(A-B)=\cos A\cos B+\sin A\sin B\) \(\cos(90-\theta)=\cos90\cos\theta+\sin90\sin\theta\) \(\cos90=0\) and \(\sin90=1\) \(\cos(90-\theta)=0\cos\theta+1\sin\theta\) \(=0+\sin\theta\) \(=\sin\theta\)

OpenStudy (anonymous):

wow thanks thats a very detailed answer and very clear

OpenStudy (ajprincess):

Welcome:)

OpenStudy (anonymous):

:)